OpenStudy (anonymous):

Can someone please help me with this quiz? Simplify: square root of 5 open parentheses 8 plus 3 square root of 6 close parentheses 8 square root of 5 end root plus 3 square root of 30 5 square root of 10 130 None of the above

OpenStudy (mathmate):

Is this what you mean? \[\sqrt{5}(8+3\sqrt{6})\]

OpenStudy (mathmate):

and the answer choices are (?): \[8\sqrt{5}+3\sqrt{30}\] \[5\sqrt{10}\] \[130\]

OpenStudy (mathmate):

plus \[none of the above\]